Calculate total no. of electrons having m = 0 in Cr atom – 

Grade:11

1 Answers

jay
6 years ago
Chromium has electronic configuration as 1s^2 2s^2 2p^6 3s^2 3p^6 4s^1 3d^5 . So since s subshell has m =0 and p subshell has m =-1 ,0,-1 and d has m=-2,-1,0,-1,-2 . So p has 1 orbital having m=0 and 1 d orbital having m =0 . So total no of electrons will be =2+2+2+2+2+1+1=12